OpenStudy (anonymous):

Suppose a is an integer. If a mod 11 = 6, what is 4a mod 11?

OpenStudy (amistre64):

what is 4*6?

OpenStudy (amistre64):

a = 6 (mod 11) 4a = 4*6 (mod 11)

OpenStudy (anonymous):

so 24 mod 11?

OpenStudy (amistre64):

yes, but its customary to reduce that to a number between from 0 to 10

OpenStudy (amistre64):

how many times does 11 go into 24, evenly?

OpenStudy (anonymous):

24 mod 11 is 3 correct?

OpenStudy (anonymous):

i mean 2

OpenStudy (amistre64):

hmm, 22 = 0 23 = 1 24 = 2 , id stick with 2 :)

OpenStudy (anonymous):

so 4a = 2? a = 2/4?

OpenStudy (anonymous):

or 1/2

OpenStudy (amistre64):

a = 6

OpenStudy (amistre64):

a = 6 (mod 11) means the same as: a mod 11 = 6

OpenStudy (amistre64):

a is any value in an equivalence class .... 6,17,28,39,51, ... 6+11n

OpenStudy (anonymous):

so 127 mod 11 = 6?

OpenStudy (anonymous):

or am i confusing this? lol

OpenStudy (amistre64):

127 - 6 = 121 = 11(11) so yes

OpenStudy (amistre64):

all a mod value is, is the remainder that is left over after division

OpenStudy (anonymous):

so 4a mod 11 would be 4(127) mod 11?

OpenStudy (amistre64):

since a = 6 (mod 11) and 127 = 6 (mod 11) then 4(127) = 4(6) (mod 11) = 2 (mod 11)

OpenStudy (anonymous):

ok its starting to make sense now, thank you so much

OpenStudy (amistre64):

youre welcome :)